OpenStudy (anonymous):

Is it possible to create a fourth-degree polynomial with only two real zeros?

OpenStudy (ash2326):

yes, that's possible. A polynomial with real coefficients if it has complex roots then they will occur as conjugate pair, suppose 2+3i is a root, it has to have 2-3i as a root as well. Question mentions that 2 roots are real, other two would be a complex conjugate pair. Certainly we can have such a case, but we can't have 1 real root and 3 complex roots for a 4th degree polynomial. Do you follow?

OpenStudy (anonymous):

I think I understand, thank you! But how would I go about making one?

OpenStudy (ash2326):

That should be easy, choose any two real values? tell me two real values

OpenStudy (anonymous):

I hate sound stupid, but those are just rational numbers, right?

OpenStudy (ash2326):

Real numbers will include rational+ irrational numbers. You can tell any two simple +ve or -ve numbers

OpenStudy (anonymous):

Does 12 and 6 work?

OpenStudy (ash2326):

yes, for sure

OpenStudy (ash2326):

ok, we have out first two zeros, other zeros would be complex. Ok the polynomial will be of the form \[(x-12)(x-6)(x-A)(x-B)\] A and B are complex conjugate roots

OpenStudy (ash2326):

Do you know complex no.s ?

OpenStudy (anonymous):

I learned about them a while ago, but I don't know. Is it with imaginary i?

OpenStudy (ash2326):

yes, you're right.

OpenStudy (anonymous):

Okay! So A and B would have imaginary numbers.

OpenStudy (ash2326):

yes, and they would be conjugate. of the form A=c+di B=c-di